# Uncovering Hidden Connections: Iterative Search and Reasoning for Video-grounded Dialog

### v3

@@ 4.4 Ablation Study
    $\hat{\mathbf{E}}_{q}=\mathbf{W}_{f}[\mathbf{E}_{q}^{txt}||\mathbf{E}_{q}^{vis}]+\mathbf{b},$ (14)
@@
  • w/o GPT-2 decoder. In this derivative, the GPT-2 decoder is supplanted with the LSTM-based auto-regressive decoder that is previously utilized in [16].
Table V presents a meticulous evaluation of the ISR model against its various derivatives using the AVSD@DSTC7 dataset. We can obtain the following salient insights from the results:
@@
4.5 Comparison with LVLMs
@@
It is well known that LVLMs have achieved excellent performance on several video-language tasks, e.g., video question answering and video captioning. Therefore, to verify their effectiveness on the video-grounded dialog task, we chose representative Video-LLaMA [41] and Video-LLaVA [57] methods to compare with ours, and the experimental results are shown in Table VI.
@@
From the displayed results, we can see that our model significantly outperforms the two LVLMs over all evaluation metrics on two datasets. For a broad comparison, we have also utilized ChatGPT for evaluation [^8] https://github.com/mbzuai-oryx/Video-ChatGPT/tree/main/quantitative_evaluation., focusing on sentence semantics eliminates structure effects. It should be noted that Accuracy metric refers to whether the predicted answer semantically matches the ground truth, while the Score is an value between 0 and 5, with 5 indicating the highest meaningful match. In this case, our method also achieves the best performance. Crucially, compared to 7B parameters of the two LVLMs, our method yields such results using only about 550M parameters. These findings underscore the challenges involved in adapting LVLMs to video-grounded dialog task and further corroborate the efficacy of our method in uncovering the complex interplays within video-grounded dialog.
4.6 Decoding Methods
@@
Table VII furnishes an in-depth analysis of various decoding strategies deployed within the domain of video-grounded dialog. A synopsis of our findings is delineated below:
@@
  • Greedy Search. While greedy search that selects the most probable subsequent word in each instance exhibits laudable performance (BLEU-1 score of 0.756), it can sometimes overlook globally coherent or intricate responses due to its short-sighted selection strategy.
@@
  • Nucleus Sampling. Posting a BLEU-1 score of 0.695, nucleus sampling amalgamates randomness by selecting from a subset of the most likely subsequent words. The introduction of diversity in responses might interfere with answer coherence in comparison to its counterparts within video-grounded dialogs.
@@
  • Contrastive Search. Achieving a commendable BLEU-1 score of 0.773, the contrastive approach evaluates divergent options, potentially aiding in discerning the nuances of video-grounded responses. Despite its superiority over the first two methods, it could not surpass beam search in our assessment.
@@
  • Beam Search. Clinching the pinnacle with a BLEU-1 score of 0.782, beam search meticulously evaluates multiple trajectories during decoding, ultimately selecting the most contextually rich sequence. Its dominance across all evaluations attests to its suitability for tasks requiring precise alignment with videos.
@@
TABLE VII: Performance comparison among the variants of different generation methods on the AVSD@DSTC7 dataset. The best results are highlighted in bold.
@@
Greedy Search 0.756 0.619 0.509 0.422 0.294 0.598 1.241
Nucleus Sampling 0.695 0.534 0.417 0.328 0.262 0.539 0.981
Contrastive Search 0.773 0.641 0.534 0.443 0.299 0.604 1.299
Beam Search 0.782 0.655 0.552 0.469 0.305 0.619 1.331
@@
Figure 4: Parameter analysis of our model in terms of the (a) the number of dialog turns, (b) the selection threshold $p$, (c) the number of iterations $I$.
4.7 Parameter Analysis
@@
4.7.1 On the number of dialog turns
@@ 4.7.3 On the number of iterations $I$
4.8 Case Study
@@
Figure 5: Two visualization examples.
5 Conclusion And Future Work

### v4

@@ 4.4. Ablation Study
  • w/ single attention. Instead of adopting four separate attention networks, we use only one attention mechanism to obtain $\mathbf{E}_{q}^{q}$, $\mathbf{E}_{q}^{v}$, $\mathbf{E}_{v}^{q}$, and $\mathbf{E}_{v}^{v}$.
  • w/o weight matrix. Remove the weight matrix obtained in Eqn. (7). This means removing the weight coefficients in Eqn. (9), treating different modalities equally.
@@
    (14) $\hat{\mathbf{E}}_{q}=\mathbf{W}_{f}[\mathbf{E}_{q}^{txt}||\mathbf{E}_{q}^{vis}]+\mathbf{b},$
@@
  • w/o GPT-2 decoder. In this derivative, the GPT-2 decoder is supplanted with the LSTM-based auto-regressive decoder that is previously utilized in (Le et al., 2019b).
Table 5 presents a meticulous evaluation of the ISR model against its various derivatives using the AVSD@DSTC7 dataset. We can obtain the following salient insights from the results:
@@
  • The performance of using a single attention mechanism, i.e., w/ single attention, is significantly worse compared to ISR model with four separate attention networks, thereby validating the rationality and validity of this setup.
  • The results are significantly lower than those of our ISR method, indicating that treating different modalities equally is inappropriate. This is reasonable, as the contributions of different modalities (e.g., question and video) are not the same.
@@
Table 6. Performance comparison with the large video-language models. The version of ChatGPT is “gpt-3.5-turbo” and the best performance is highlighted in bold.
Methods Objective Metric ChatGPT
BLEU-1 BLEU-4 METEOR ROUGE-L CIDEr Accuracy Score
AVSD@DSTC7
Zero-shot Setting
Video-LLaMA-7B 0.360 0.115 0.210 0.367 0.383 28.533 2.551
Video-LLaVA-7B 0.422 0.142 0.239 0.418 0.447 33.158 2.918
Fine-tuning Setting
Video-LLaVA-7B 0.749 0.430 0.287 0.580 1.199 50.318 3.477
ISR 0.792 0.472 0.315 0.623 1.344 52.783 3.645
AVSD@DSTC8
Zero-shot Setting
Video-LLaMA-7B 0.345 0.110 0.201 0.364 0.444 28.737 2.575
Video-LLaVA-7B 0.410 0.137 0.228 0.416 0.527 33.275 2.998
Fine-tuning Setting
Video-LLaVA-7B 0.737 0.410 0.278 0.569 1.142 52.336 3.684
ISR 0.767 0.464 0.304 0.610 1.288 51.812 3.658
4.5. Comparison with LVLMs
@@
It is well known that LVLMs have achieved excellent performance on several video-language tasks, e.g., video question answering and video captioning. Therefore, to verify their effectiveness on the video-grounded dialog task, we chose representative Video-LLaMA (Zhang et al., 2023b) and Video-LLaVA (Lin et al., 2023) methods to compare with ours, and the experimental results are shown in Table 6.
@@
From the displayed results, under both zero-shot and fine-tuning settings, it is evident that our model significantly outperforms the two LVLMs across all objective metrics on both datasets. For a broad comparison, we have also utilized ChatGPT for evaluation [^8] https://github.com/mbzuai-oryx/Video-ChatGPT/tree/main/quantitative_evaluation., focusing on sentence semantics eliminates structure effects. The structured prompt is shown in Fig. LABEL:fig:prompt. It should be noted that Accuracy metric refers to whether the predicted answer semantically matches the ground truth, while the Score is an value between 0 and 5, with 5 indicating the highest meaningful match. In this case, our ISR model similarly beats the fine-tuned Video-LLaVA-7B model on the AVSD@DSTC7 dataset. On the AVSD@DSTC8 dataset, our approach is inferior to the fine-tuned Video-LLaVA-7B model, perhaps because the rich knowledge embedded in LVLM comes into play in the face of more challenging data.
Crucially, compared to 7B parameters of the two LVLMs, our method yields such results using only about 550M parameters (See Table 10). These findings underscore the challenges involved in adapting LVLMs to video-grounded dialog task and further corroborate the efficacy of our method in uncovering the complex interplays within video-grounded dialog.
4.6. Effect of Different Encoders/Decoders
In Table 7, we designed several variants to investigate the effects of different encoder (e.g., Faster R-CNN, ResNext, and I3D) and decoder (e.g., GPT-2-Base, LSTM, GPT-2-Large) configurations. From the displayed results, we can observe that modifying the encoders and decoders does not significantly impact the overall model performance, resulting only in minor fluctuations. And in these results, the performance of the variant with Faster R-CNN as encoder and GPT-2-Large as decoder is optimal. Besides, we also found that the variant using lower-quality I3D features could still achieve performance comparable to the optimal baseline (DialogMCF (Chen et al., 2023)). These results fully validate the robustness and generalization of our proposed method.
Table 7. Performance comparison among the variants with different encoders and decoders on the AVSD@DSTC7 dataset.
Encoder Decoder BLEU-1 BLEU-4 METEOR ROUGE-L CIDEr
Faster R-CNN GPT-2-Base 0.792 0.472 0.315 0.623 1.344
ResNext GPT-2-Base 0.785 0.463 0.310 0.619 1.326
I3D GPT-2-Base 0.780 0.459 0.308 0.616 1.317
Faster R-CNN LSTM 0.774 0.447 0.298 0.609 1.276
Faster R-CNN GPT-2-Large 0.805 0.481 0.322 0.634 1.382
4.7. Effect of Different Inputs
In order to explore the impact of the different inputs of generator, we eliminated them individually, as shown in Table 8. This will result in three variants (with the question not being removed): w/o video, w/o caption, and w/o history. From the results, we can observe that removing any input leads to a significant decline in model performance. Moreover, the impact of removing different inputs varies, with the influence decreasing in the following order: caption, dialog history, and video. This phenomenon has also been validated in related studies (Chen et al., 2023; Le et al., 2019b).
Table 8. Performance comparison among the variants with different inputs of generator on the AVSD@DSTC7 dataset. The best results are highlighted in bold.
Variant BLEU-1 BLEU-4 METEOR ROUGE-L CIDEr
ISR 0.792 0.472 0.315 0.623 1.344
w/o video 0.775 0.456 0.307 0.612 1.316
w/o caption 0.757 0.439 0.293 0.603 1.288
w/o history 0.768 0.441 0.302 0.607 1.301
4.8. Decoding Methods
@@
Table 9 furnishes an in-depth analysis of various decoding strategies deployed within the domain of video-grounded dialog. A synopsis of our findings is delineated below:
@@
  • Greedy Search. While greedy search that selects the most probable subsequent word in each instance exhibits laudable performance (BLEU-1 score of 0.765), it can sometimes overlook globally coherent or intricate responses due to its short-sighted selection strategy.
@@
  • Nucleus Sampling. Posting a BLEU-1 score of 0.702, nucleus sampling amalgamates randomness by selecting from a subset of the most likely subsequent words. The introduction of diversity in responses might interfere with answer coherence in comparison to its counterparts within video-grounded dialogs.
@@
  • Contrastive Search. Achieving a commendable BLEU-1 score of 0.780, the contrastive approach evaluates divergent options, potentially aiding in discerning the nuances of video-grounded responses. Despite its superiority over the first two methods, it could not surpass beam search in our assessment.
@@
  • Beam Search. Clinching the pinnacle with a BLEU-1 score of 0.792, beam search meticulously evaluates multiple trajectories during decoding, ultimately selecting the most contextually rich sequence. Its dominance across all evaluations attests to its suitability for tasks requiring precise alignment with videos.
@@
Table 9. Performance comparison among the variants of different generation methods on the AVSD@DSTC7 dataset. The best results are highlighted in bold.
@@
Greedy Search 0.765 0.628 0.514 0.422 0.300 0.605 1.254
Nucleus Sampling 0.702 0.544 0.422 0.328 0.268 0.545 0.994
Contrastive Search 0.780 0.649 0.545 0.447 0.307 0.606 1.312
Beam Search 0.792 0.665 0.560 0.472 0.312 0.623 1.344
@@
Figure 5. Parameter analysis of our model in terms of (a) the number of dialog turns, (b) the selection threshold $p$, (c) the number of iterations $I$.
4.9. Parameter Analysis
@@
4.9.1. On the number of dialog turns
@@ 4.9.3. On the number of iterations $I$
Table 10. Comparison on the number of trainable parameters, memory usage, and inference speed.
Methods Parameter Memory Speed
MTN (Le et al., 2019b) 289.11M 2.59G 3.42s
BiST (Le et al., 2020) 17.99M 8.08G 1.31s
RLM (Li et al., 2021b) 523.75M 2.27G 1.24s
ISR 546.09M 2.31G 1.27s
4.10. Efficiency Analysis
In order to more fully assess the efficiency of our model, we have calculated the number of trainable parameters, memory usage, and inference speed of our model as well as the representative baselines, which are presented in Table 10. The entire computation is done on AVSD@DSTC7 dataset using a NVIDIA-V100-32G GPU. It is worth noting that since existing work does not emphasize the above factors, we only have compared open-source methods. As reported in Table 10, in comparison to the similar method RLM (Li et al., 2021b), our method ISR achieves better performance without significantly increasing memory usage and inference speed.
Figure 6. Visualization of question-video weights.
(a) Case A.
(b) Case B.
Figure 7. Two visualization examples.
4.11. Visualization of Question-Video Weights
To visually present the validity of visual content comprehension, we provided a case and visualized the question-video weights for demonstration, as shown in Figure 6. In this case, the question “does he put the box back on the counter as well ?” is represented along the Y-axis. For the video, three sampled frames are displayed due to space constraints. Each frame contains the bounding boxes of six objects (fixed setting) along with their corresponding categories, which are mapped to the X-axis. Each grid cell represents the weight between the embedding of a given word in the question and the corresponding object feature in the video frame. From the case, we can observe that word embedding and object feature corresponding to the same concept, such as “counter” and “box” highlighted in red boxes, exhibit similarity. This provides convincing evidence to demonstrate the validity of the proposed iterative reasoning network.
4.12. Case Study
@@
5. Conclusion and Future Work
@@
Acknowledgements.
This work is supported in part by the National Natural Science Foundation of China, No.:62376140 and No.:U23A20315; Shenzhen College Stability Support Plan, No.:GXWD20220817144428005; the Science and Technology Innovation Program for Distinguished Young Scholars of Shandong Province Higher Education Institutions, No.: 2023KJ128; Pengcheng Laboratory Major Research Projects, No.:PCL2023A08; and in part by the Special Fund for Taishan Scholar Project of Shandong Province.